# Spatial-frequency Dual-Domain Feature Fusion Network for Low-Light Remote Sensing Image Enhancement

### v1

@@ I Introduction
Although deep convolutions can expand the receptive field, as the network depth increases, the loss of detailed information gradually occurs, making it challenging for models to establish long correlations in RS images. Recently, a number of transformer methods[17, 18, 19, 20, 21], with Vision Transformer[22] as a prominent example, have successfully addressed the challenge of establishing long-range dependencies by implementation of self-attention mechanisms. However, the quadratic complexity problem remains difficult to solve. Additionally, these methods tend to overfit when applied to small datasets, making practical application challenging. Fig. 1 (a) presents visual results and number of parameters for several representative methods. It can be observed that SNR, built upon the transformer, achieves suboptimal visual results, yet concurrently bears an onerous computational burden.
@@
Most existing methodss[23, 24, 25] process images in the spatial domain based on a pixel perspective. In contrast, we shift our focus to the Fourier domain, motivated by two unique phenomena inherent in the Fourier domain: 1) Under Fourier transformation, global information of the image can be extracted without introducing a large number of parameters[26]. 2) It is well-known that Fourier phase preserves high-level semantics, while amplitude contains low-level features[27]. Fig. 2 displays the visualization image of exchanging the Fourier amplitude and phase between normal and low-light images. It can be observed that brightness degradation exists in the amplitude, while other information is preserved in the phase. This phenomenon suggests that the Fourier transform can decouple the degradation of low-light images, breaking down the complex coupling problem into two relatively easier-to-solve sub-problems. Existing methods have explored the frequency domain information, but still present many shortcomings. For example, Wang et al.[26] opts to simultaneously recover both amplitude and phase information without considering spatial information. Li et al.[28] only performs Fourier decomposition on images within the network, yet the loss function continues to make comparisons with the ground truth (GT) solely in the spatial domain, lacking effective supervision. This undoubtedly makes the recovery process more difficult. Therefore, the challenge of how to integrate frequency domain information with spatial features and efficiently apply this integration for the enhancement of remote sensing low-light images remains unresolved.
@@
  • We design an Information Affine Fusion Module to facilitate information interaction across scales, stages, and domains within the two-stage network. This module promotes information fusion between the two stages and enhances the representation of global context.
@@ II-A Low-light image enhancement
Figure 4: The specific structure of the proposed IFAM consists of two key components: the Information Fusion Module and the Information Mapping Module.
II-B Fourier Transform in Neural Networks
Recently, due to the unique properties of images in the Fourier space, such as capturing global representations, Fourier frequency information [26, 28] has attracted widespread attention. A series of works have been based on Fourier domain transformations to enhance the generalization performance of networks. For instance, Mao et al. [46] designed a residual Fourier block to address issues like the potential neglect of low-frequency information and the difficulty in modeling long-distance information in Resblocks. Roman et al. [47] used fast Fourier convolution to design a large mask inpainting network, achieving impressive results even in challenging scenarios. Dario et al. [48] employed Fourier loss to aid networks in recovering high-frequency information in super-resolution images. While these works explore the application of Fourier information in various scenarios, few have considered the relationship between frequency domain information and remote sensing images. In contrast, our proposes DFFN effectively decouples dark light degradation, achieving excellent enhancement results.
@@ III-C2 Information Affine Module
SCI[50] CVPR’22 14.78 0.525 0.238 12.14 0.194 0.901 13.08 0.205 0.745
SNR[13] CVPR’22 21.22 0.834 0.097 13.31 0.246 0.722 23.17 0.760 0.179
LLFormer[51] AAAI’23 23.25 0.819 0.105 16.66 0.376 0.658 23.11 0.725 0.221
CUE[52] ICCV’23 21.68 0.769 0.151 15.70 0.343 0.717 21.47 0.682 0.263
FourLLIE[26] ACMM’23 20.02 0.818 0.088 14.49 0.350 0.579 19.78 0.643 0.309
LANet[53] IJCV’23 21.74 0.816 0.101 11.88 0.089 0.804 18.72 0.695 0.231
NeRCo[54] CVPR’23 22.94 0.783 0.102 16.73 0.513 0.381 18.79 0.604 0.326
@@
where $\otimes$ indicates the convolution operation.
Figure 6: Distribution analysis of the iSAID-dark dataset. (a) T-SNE map indicates that the proposed dataset exhibits a broader data range. (b) demonstrates the brightness range of normal-light (blue) and low-light images (green). (c) The whole dataset covers a wide range of SNR distribution ranging from 1 to 25 and the SNR values center at the range of [0,10], indicating the challenging noise levels.
@@ III-D Loss Function
where $\left\|\cdot\right\|_{1}$ presents the Mean Absolute Error(MAE), $\alpha$ and $\beta$ are the weight factor, and we empirically set them to 0.05. Therefore, the overall loss is :
@@
Due to the particularity of remote sensing images, collecting paired normal/low-light image sets is almost an impossible task. In this work, inspired by the significant achievements of previous synthetic datasets in various low-quality image enhancement tasks, we introduced a synthesis mechanism to generate the corresponding low-light remote sensing dataset. Additionally, to validate the reliability of the proposed DDFN and iSAID-dark in real low-light environments, we collected real nighttime remote sensing images by capturing different scenes with drones at night, forming a test dataset named dark remote sensing (darkrs), which comprises 86 test images. The two proposed datasets are shown in Fig. 5.
@@ IV-A Data Collection
The iSAID[55] dataset has collected a large number of high-resolution aerial images from various scenes including rural and urban areas. Therefore, we selected 751 images from the iSAID test set as our benchmark data. Considering the significant differences in size and shape of remote sensing images, and the difficulty of training high-resolution images in mainstream networks, and in order to increase the diversity of scenes included in the dataset, we performed multiple random crops on these 751 images. The cropped images are resized to 500x500. Subsequently, we divided them into 3755 training image pairs and 66 validation image pairs.
TABLE II: Comparison with state-of-the-art methods on the no-reference dataset, with the best results indicated in bold and the second-best results underlined.
@@
SCI[50] CVPR’22 3.968 3.657 3.912 4.165 5.450 4.230 0.0004
SNR[13] CVPR’22 4.125 4.458 5.503 5.844 3.456 4.677 40.0842
LLFormer[51] AAAI’23 3.960 3.890 3.629 4.175 4.502 4.031 24.5183
CUE[52] ICCV’23 3.989 3.722 3.977 4.068 3.708 3.892 0.2407
FourLLIE[26] ACMM’23 3.941 3.685 3.550 4.056 4.119 3.870 1.4855
LANet[53] IJCV’23 4.322 3.992 3.828 4.493 3.478 4.022 0.0422
NeRCo[54] CVPR’23 4.092 4.052 4.048 4.007 4.140 4.067 7.8414
@@ IV-C Data Distribution
Fig. 6 (a) displays the T-SNE visualization results of our proposed iSAID-dark dataset compared to other datasets, demonstrating that iSAID-dark encompasses common features of LOL and SICE, indicating a broader data range in our dataset. Fig. 6 (b) illustrates the brightness distribution of our proposed dataset. We also plotted the SNR distribution of the dataset in Fig. 6 (c) to showcase the noise level of the images. The SNR distribution indicates that our dataset possesses a wide and challenging SNR range.
@@ V-A3 Implementation Details
Our model is trained on the LOL dataset and the iSAID-dark dataset separately. To facilitate better convergence of the model, we set the initial learning rate to 0.0002, and then linearly decrease it by half every 100 epochs, with a total of 500 epochs for training. We use the ADAM optimizer for model optimization. All experiments are conducted on a Windows computer equipped with an NVIDIA RTX 3090 (24GB RAM) for training and testing. Additionally, we set the batch size to 12, randomly crop each image into 256$\times$256 patches, and augment these patches using data augmentation techniques such as horizontal and vertical flips.
@@ V-B1 Quantitative Results
We first quantitatively compare DFFN with state-of-the-art methods. Table I presents the PSNR, SSIM, and LPIPS results on reference-based datasets. From Table I, it can be seen that our method achieved the best or second-best results on both reference datasets. It is worth noting that the results shown for iSAID-dark are obtained by testing with weights trained on LOL, indicating that models trained on existing datasets may not be suitable for low-light remote sensing image enhancement.
Table II presents the NIQE results on the no-reference dataset. It can be observed from Table II that our method achieves the best on all datasets, which fully demonstrates the superiority of DFFN. We also present the parameter count of the models in Table II. The results indicate that although our method cannot compete with no-reference methods in terms of complexity, the performance of no-reference methods on low-light remote sensing datasets is extremely unsatisfactory (for example, SSIM of Zero-DCE++ on iSAID-dark is only 0.055). In contrast, our method achieves a remarkable balance between model complexity and performance.
@@ V-B2 Qualitative Results
Meanwhile, Fig. 12 displays the NIQE distributions of the comparison methods across five test datasets. It is evident that DFFN exhibits a more concentrated data distribution, which strongly indicates the superior generalization capability of our model across various scenarios. Additionally, when all test set data is combined and plotted together (denoted as ALL), DFFN’s data appears more centralized. This further demonstrates the exceptional stability of our model.
TABLE V: The ablation study for using transformer blocks, ’One_block’ indicates that each block uses only one self-attention block and one FFN, while ’Stack_block’ signifies that multiple self-attention blocks and FFNs are stacked within each transformer block.
@@ V-C Ablation Study
In this subsection, we conducted extensive ablation experiments to validate the effectiveness of the proposed modules, the two-stage network, and the fusion of spatial and frequency domains.
@@ V-C2 Contribution of Dual-Domain Fusion
In this subsection, we discuss the effectiveness of integrating spatial and frequency domain information and the effectiveness of different integration methods, while also exploring the effectiveness of the proposed IFAM. $M_{1}$ denotes the use of spatial information only in a two-level network. $M_{2}$ and $M_{3}$ adopt a sequential method to learn spatial and frequency domain information, respectively. Specifically, $M_{2}$ first learns spatial information (as shown in Fig. 13), while $M_{3}$ first learns frequency domain information. $M_{4}$ has the same training architecture as ours, but none of the four configurations use IFAM.
As shown in Table IV, the model $M_{1}$ utilizing only spatial information performs poorly, indicating the effectiveness of integrating spatial-frequency dual-domain information. Compared to $M_{2}$ and $M_{3}$ ,$M_{4}$, which learns in parallel, achieves the best performance in most metrics. This suggests that compared to the parallel approach, serial learning is more prone to losing information from one domain while learning information from another. By employing parallel computation, this information loss is effectively mitigated.
@@
V-C3 Compared with transformer blocks

### v2

@@ I Introduction
Although deep convolutions can expand the receptive field, as the network depth increases, the loss of detailed information gradually occurs, making it challenging for models to establish long correlations in RS images. Recently, a number of transformer methods[17, 18, 19, 20, 21], with Vision Transformer[22] as a prominent example, have successfully addressed the challenge of establishing long-range dependencies by implementation of self-attention mechanisms. However, the quadratic complexity problem remains difficult to solve. Additionally, these methods tend to overfit when trained on small datasets, significantly reducing their reliability in practical applications. Fig. 1 (a) presents visual results and number of parameters for several representative methods. It can be observed that SNR, built upon the transformer, achieves suboptimal visual results, yet concurrently bears an onerous computational burden.
@@
Most existing methods[23, 24, 25] process images in the spatial domain based on pixel perspectives. For instance, Xing et al.[26] proposed an unsupervised GAN framework for low-light remote sensing image enhancement based on the self-similarity characteristics of remote sensing images. Singh et al.[27] designed a multi-resolution Rnet that preserves high-resolution information while retaining low-resolution standards through different data streams. In contrast, we shift our focus to the Fourier domain, motivated by two unique phenomena inherent in the Fourier domain: 1) Under Fourier transformation, global information of the image can be extracted without introducing a large number of parameters[28]. 2) It is well-known that Fourier phase preserves high-level semantics, while amplitude contains low-level features[29]. Fig. 2 displays the visualization image of exchanging the Fourier amplitude and phase between normal and low-light images. It can be observed that brightness degradation exists in the amplitude, while other information is preserved in the phase. This phenomenon suggests that the Fourier transform can decouple the degradation of low-light images, breaking down the complex coupling problem into two relatively easier-to-solve sub-problems. Existing methods have explored the frequency domain information, but still present many shortcomings. For example, Wang et al.[28] opts to simultaneously recover both amplitude and phase information without considering spatial information. Li et al.[30] only performs Fourier decomposition on images within the network, yet the loss function continues to make comparisons with the ground truth (GT) solely in the spatial domain, lacking effective supervision. This undoubtedly makes the recovery process more difficult. Therefore, the challenge of how to integrate frequency domain information with spatial features and efficiently apply this integration for the enhancement of remote sensing low-light images remains unresolved.
Additionally, multi-stage architecture networks[31, 4, 28, 32] often face issues with insufficient information interaction. Most existing works directly use the output of the previous stage as the input for the next stage. However, such methods do not consider the rich information contained in the intermediate features, which can easily lead to information loss. Other works adopt a one-to-one correspondence approach to pass information from the previous stage to the next stage, effectively compensating for the information loss and detail blurring caused by the upsampling and downsampling processes. However, these solutions still only consider the interaction of information at the same scale, neglecting the interaction of cross-scale features within and between stages, making it difficult to obtain more refined contextual features.
@@
  • We design an Information Fusion Affine Module to facilitate information interaction across scales, stages, and domains within the two-stage network. This module promotes information fusion between the two stages and enhances the representation of global context.
@@ II-A Low-light image enhancement
Figure 4: The specific structure of the proposed IFAM consists of two key components: the Information Fusion Module and the Information Mapping Module. $U_{i}$ represents the feature obtained after the $i$-th upsampling in the Phase refinement stage.
II-B Multi-stage Network
The multi-stage network is designed to decompose challenging tasks into several easier sub-tasks and solve them progressively to achieve optimal results. However, current multi-stage networks still have some issues. Specifically, existing multi-stage networks can be divided into the following two categories: 1) ”result fusion” and 2) ”cross-level fusion”.
The first category often uses the output of the previous stage as the input for the next stage. For example, Zhu et al. [31] proposed an Edge-Enhanced Multi-Exposure Fusion Network (EEMEFN), where they initially brighten dark images and then enhance the edges of the preliminary enhanced results to obtain the final restored outcome. Li et al. [4] utilized a recurrent network to learn low-light image enhancement, with each stage taking the attention map and the previous stage’s output as input, then performing recursive learning through the same residual module. Wang et al. [28], in their proposed dual-stage network, first enhanced image brightness by estimating magnitude information, followed by the introduction of a signal-to-noise ratio (SNR) map to restore image details. Although these methods achieve good visual effects, they all neglect the rich intermediate features, and solely using the preliminarily enhanced image can lead to information loss, resulting in blurred details.
The second category aims to integrate rich contextual features across multiple stages, alleviating information loss and gradient vanishing issues in deep networks. For instance, in the proposed Deep Generalized Unfolding Network (DGUNet), Mou et al. [32] passed the decoder features from each stage to the corresponding module in the next stage to capture rich informational features. However, these methods typically only consider the interactions between paired features and do not account for feature interactions across different scales.
In contrast, our proposed IFAM effectively combines feature information across different stages and scales. By utilizing dynamic fusion and affine filtering for adaptive learning, it significantly enhances the network’s representational capacity.
II-C Fourier Transform in Neural Networks
Recently, due to the unique properties of images in the Fourier space, such as capturing global representations, Fourier frequency information [28, 30, 50] has attracted widespread attention. A series of works have been based on Fourier domain transformations to enhance the generalization performance of networks. For instance, Mao et al. [51] designed a residual Fourier block to address issues like the potential neglect of low-frequency information and the difficulty in modeling long-distance information in Resblocks. Roman et al. [52] used fast Fourier convolution to design a large mask inpainting network, achieving impressive results even in challenging scenarios. Dario et al. [53] employed Fourier loss to aid networks in recovering high-frequency information in super-resolution images. Meanwhile, the field of remote sensing has seen extensive exploration of frequency domain information [54]. For example, Zhou et al. [55] demonstrated excellent performance in the pan-sharpening task by exploring the relationship between Pan images and low-resolution multispectral (MS) images with frequency domain information. Yu et al. [56] proposed a spiking neural network in the frequency domain (SNNFD), which enhances the model’s transferability and feature extraction capability for buildings of different sizes by integrating frequency and spatial domain learning. Inspired by these works, we proposed DFFN based on Fourier transform, which effectively decouples low-light degradation and achieves superior enhancement results.
@@ III-B2 Dual-Domain Phase Block
Figure 6: Distribution analysis of the iSAID-dark dataset. (a) T-SNE map indicates that the proposed dataset exhibits a broader data range. (b) demonstrates the brightness range of normal-light (blue) and low-light images (green). (c) In this figure, the range distribution of SNR values for both the LOL dataset and the proposed dataset is shown. It can be observed that the proposed dataset has a lower SNR range, indicating that its noise levels are more challenging.
@@ III-C2 Information Affine Module
SCI[58] CVPR’22 14.78 0.525 0.238 12.14 0.194 0.901 13.08 0.205 0.745 13.73 0.168 0.854
SNR[13] CVPR’22 21.22 0.834 0.097 13.31 0.246 0.722 23.17 0.760 0.179 22.06 0.711 0.238
LLFormer[59] AAAI’23 23.25 0.819 0.105 16.66 0.376 0.658 23.11 0.725 0.221 - - -
CUE[60] ICCV’23 21.68 0.769 0.151 15.70 0.343 0.717 21.47 0.682 0.263 19.16 0.631 0.349
FourLLIE[28] ACMM’23 20.02 0.818 0.088 14.49 0.350 0.579 19.78 0.643 0.309 17.58 0.503 0.302
LANet[61] IJCV’23 21.74 0.816 0.101 11.88 0.089 0.804 18.72 0.695 0.231 21.95 0.690 0.247
NeRCo[62] CVPR’23 22.94 0.783 0.102 16.73 0.513 0.381 18.79 0.604 0.326 20.94 0.568 0.377
@@
where $\otimes$ indicates the element-wise multiplication.
@@ III-D Loss Function
where $\left\|\cdot\right\|_{1}$ presents the Mean Absolute Error (MAE), $\alpha$, $\beta$ and $\gamma$ are the weight factor, and we empirically set them to 0.05. Therefore, the overall loss is :
@@
Due to the particularity of remote sensing images, collecting paired normal/low-light image sets is almost an impossible task. In this work, inspired by the significant achievements of previous synthetic datasets in various low-quality image enhancement tasks, we introduced a synthesis mechanism to generate the corresponding low-light remote sensing dataset. Additionally, to validate the reliability of the proposed DFFN and iSAID-dark in real low-light environments, we collected real nighttime remote sensing images by capturing different scenes with drones at night, forming a test dataset named dark remote sensing (darkrs), which comprises 86 test images. The two proposed datasets are shown in Fig. 5.
@@ IV-A Data Collection
The iSAID[63] dataset has collected a large number of high-resolution aerial images from various scenes including rural and urban areas. Therefore, we selected 751 images from the iSAID test set as our benchmark data. Considering the significant differences in size and shape of remote sensing images, and the difficulty of training high-resolution images in mainstream networks, and in order to increase the diversity of scenes included in the dataset, we performed multiple random crops on these 751 images. The cropped images are resized to 500x500. Subsequently, we divided them into 3755 training image pairs and 66 validation image pairs. To further validate the recovery performance at high resolutions, we constructed a corresponding high-resolution dataset called iSAID-dark (high-pixel), which includes 72 test images, all with a resolution of 1080p.
TABLE II: Comparison with state-of-the-art LLIE methods on the no-reference datasets, along with the number of parameters, GFLOPS, and FPS on images of 1080p resolution. The best results are indicated in bold, and the second-best results are underlined.
@@
SCI[58] CVPR’22 3.968 3.657 3.912 4.165 5.450 4.230 293 1.1 0.0004
SNR[13] CVPR’22 4.125 4.458 5.503 5.844 3.456 4.677 3 880 40.0842
LLFormer[59] AAAI’23 3.960 3.890 3.629 4.175 4.502 4.031 - - 24.5183
CUE[60] ICCV’23 3.989 3.722 3.977 4.068 3.708 3.892 3 198 0.2407
FourLLIE[28] ACMM’23 3.941 3.685 3.550 4.056 4.119 3.870 6 515 1.4855
LANet[61] IJCV’23 4.322 3.992 3.828 4.493 3.478 4.022 15 317 0.0422
NeRCo[62] CVPR’23 4.092 4.052 4.048 4.007 4.140 4.067 3 1341 7.8414
@@ IV-C Data Distribution
Fig. 6 (a) displays the T-SNE visualization results of our proposed iSAID-dark dataset compared to other datasets, demonstrating that iSAID-dark encompasses common features of LOL and SICE, indicating a broader data range in our dataset. Fig. 6 (b) illustrates the brightness distribution of our proposed dataset. Meanwhile, we plotted the SNR distribution of the proposed dataset and the LOL dataset, as shown in Fig. 6 (c), to demonstrate the noise levels in the images. The SNR distribution indicates that our dataset possesses a wide and challenging SNR range.
@@ V-A1 Datasets and Evaluation Metrics
Figure 10: The visual results on the 1080p image. The selected area is magnified to show details.
@@
Figure 13: Qualitative enhancement results of three low-light images of the same scene. From left to right, the brightness of the images gradually decreases, and the noise concentration gradually increases.
@@ V-A3 Implementation Details
Our model is trained on the LOL dataset and the iSAID-dark dataset separately. To facilitate better convergence of the model, we set the initial learning rate to 0.0002, and then linearly decrease it by half every 100 epochs, with a total of 500 epochs for training. We use the ADAM optimizer for model optimization. All experiments are conducted on a Windows computer equipped with an NVIDIA RTX 3090 (24GB RAM) for training and testing. Additionally, we set the batch size to 12, randomly crop each image into 256$\times$256 patches, and augment these patches using data augmentation techniques such as horizontal and vertical flips. For image features, we set C to 20, and subsequently, the feature levels are transformed to 40, 80, and 160.
@@ V-B1 Quantitative Results
We first quantitatively compare DFFN with state-of-the-art methods. Table I presents the PSNR, SSIM, and LPIPS results on reference-based datasets. From Table I, it can be seen that our method achieved the best results on both reference datasets. It is worth noting that the results shown for iSAID-dark are obtained by testing with weights trained on LOL, indicating that models trained on existing datasets may not be suitable for low-light remote sensing image enhancement.
Table II presents the NIQE results on the no-reference dataset. It can be observed from Table II that our method achieves the best on all datasets, which fully demonstrates the superiority of DFFN. We also present the parameter count of the models in Table . The results indicate that although our method cannot compete with no-reference methods in terms of complexity, the performance of no-reference methods on low-light remote sensing datasets is extremely unsatisfactory (for example, SSIM of Zero-DCE++ on iSAID-dark is only 0.055). In contrast, our method achieves a remarkable balance between model complexity and performance.
@@ V-B2 Qualitative Results
As shown in Fig. 10, for high-resolution images, Zero-DCE++, SCI and FourLLIE struggle to restore the image brightness and fail to address noise issues. CUE, LANet, SNR, and NeRCo exhibit unexplained artifacts or neglect color information. LLFormer, due to its large computational load, cannot process images of this resolution on our equipment. In contrast, our method maintains outstanding performance on high-resolution image.
@@
Meanwhile, we tested the recovery ability of the method in different lighting and noise environments, as shown in Fig. 13. Zero-DCE++ and SCI exhibit significantly different recovery capabilities under varying illumination conditions, and their noise repair ability is poor. LANet and FourLLIE struggle to capture color information in extremely dark environments. Due to excessive noise interference, the detail recovery capabilities of CUE and NeRCo are greatly affected. Although LLFormer and SNR, based on the Transformer architecture, demonstrate relatively excellent hierarchical correction capabilities, their performance is still unsatisfactory under excessively dark and noisy conditions. In contrast, our method achieves excellent recovery results in all three levels of adverse environments.
@@
Fig. 14 displays the NIQE distributions of the comparison methods across five test datasets. It is evident that DFFN exhibits a more concentrated data distribution, which strongly indicates the superior generalization capability of our model across various scenarios. Additionally, when all test set data is combined and plotted together (denoted as ALL), DFFN’s data appears more centralized. This further demonstrates the exceptional stability of our model.
TABLE V: The ablation study for networks with different depths. C = 10 means the initial feature channels are 10, which then expand sequentially to 2C, 4C, and 8C within the network.
[…]
C = 20 25.30 0.784 0.151 4 501 2.5765
C = 30 24.90 0.735 0.176 2 1121 5.7883
TABLE VI: The ablation study for using transformer blocks, Single Transformer Block (STB) indicates that each block uses only one self-attention block and one FFN, while Multiple Transformer Block (MTB) signifies that multiple self-attention blocks and FFNs are stacked within each transformer block.
@@ V-C Ablation Study
In this subsection, we conducted extensive ablation experiments to validate the effectiveness of the proposed modules, the two-stage network, and the fusion of spatial and frequency domains. Subsequently, we evaluated the performance of networks with different depths and the use of multi-level loss. Finally, we replaced the frequency domain block with the state-of-the-art transformer block to validate the effectiveness of the proposed model.
@@ V-C2 Contribution of Dual-Domain Fusion
In this subsection, we discuss the effectiveness of integrating spatial and frequency domain information and the effectiveness of different integration methods, while also exploring the effectiveness of the proposed IFAM and DDAB/DDPB. $M_{1}$ denotes the use of spatial information only in a two-level network. $M_{2}$ and $M_{3}$ adopt a sequential method to learn spatial and frequency domain information, respectively. Specifically, $M_{2}$ first learns spatial information (as shown in Fig. 15), while $M_{3}$ first learns frequency domain information. $M_{4}$ has the same training architecture as ours, but none of the four configurations use IFAM. $M_{5}$ indicates the absence of the proposed DDAB/DDPB.
As shown in Table IV, the model $M_{1}$ utilizing only spatial information performs poorly, indicating the effectiveness of integrating spatial-frequency dual-domain information. Compared to $M_{2}$ and $M_{3}$ ,$M_{4}$, which learns in parallel, achieves the best performance in most metrics. This suggests that compared to the parallel approach, serial learning is more prone to losing information from one domain while learning information from another. By employing parallel computation, this information loss is effectively mitigated. When DDAB/DDPB is not employed, the model’s performance declines significantly, highlighting the superiority of the proposed DDAB/DDPB.
@@
V-C3 Contribution of different depths
In this subsection, we investigate the impact of network depth on performance. Specifically, we set the number of channels expanded after the first convolutional layer to C and experiment with C = 10, 20, 30. For these three configurations, we train the network for 1000 epochs. As shown in Fig. 16, the left graph illustrates the loss decline curve on the validation set, while the right graph shows the PSNR improvement curve.
When C = 10, although the network has fewer parameters and faster inference speed, it is too shallow, resulting in slow fitting speed and poor learning quality. When C = 30, the network is too deep with more parameters, leading to overfitting and poor performance. Additionally, the network has an excessive number of parameters when C = 30, as shown in Table V. Comparatively, the network achieves optimal performance when C = 20.
Figure 16: Ablation Study on different depths. C represents the initial number of channels expanded after entering the network, and it can be seen from the figure that the network performs best when C = 20.
V-C4 Contribution of multistage loss
In this subsection, we validated the benefits of multi-level loss for the proposed model. As shown in Fig. 17, the left graph depicts the loss decline curves when using different losses, and the right graph illustrates the PSNR increase curves. ”W/o frequency” indicates the absence of all frequency domain losses, while ”W/o pixelloss” represents the absence of spatial domain loss. The remaining configurations demonstrate the results when each of the three frequency domain losses is individually omitted. From the enlarged area of the left graph, it can be observed that when only spatial or frequency domain information is used for supervision, the fitting speed is significantly slower. When both spatial and frequency domain information are used for supervision simultaneously, the convergence speed of the model is similar initially, but after 100 epochs, the model with the complete loss converges significantly faster. Notably, in the case of ”W/o fftloss,” although the model uses the corresponding frequency domain information for supervision in both stages, the lack of overall frequency domain guidance makes it more difficult for the model to fit, and the final performance is worse.
V-C5 Compared with transformer blocks